# Linear Bandits with Partially Observable Features

### v2

@@ E.2 Proof of Theorem 2
$\leq\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\right)\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)\right\|_{\infty}+\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\boldsymbol{\mu}_{\star}\right)\mathbf{\widetilde{x}}_{a}\right\|_{\infty}.$
With probability at least $1-\delta$, the event $\mathcal{M}_{\tau}$ happens for all $\tau\geq 1$ and we obtain a pair of matching sample $\widetilde{a}_{\tau}$ and $a_{\tau}$. Thus, the second term is equal to,
@@
Because $\|\mathbf{v}\|_{\infty}=\max_{i\in[d]}|{e_{i}^{\top}\mathbf{v}}|$ for any $\mathbf{v}\in\mathbb{R}^{d}$,
@@
$\leq\max_{a\in[K]}\sigma\sqrt{2\mathbf{e}_{a}^{\top}\left(\mathbf{V}_{t}\right)\mathbf{e}_{a}\log\frac{2Kt^{2}}{\delta}}$
@@
Let $\mathbf{A}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\mathbb{I}(\widetilde{a}_{\tau}=a)\phi_{a,\tau}^{-1}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$. Then the first term,
@@
Because $\check{\boldsymbol{\mu}}_{t}^{L}$ is a minimizer of Eq. 8, by Lemma 3 and Eq. 20,
$\left\|\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right\|_{\mathbf{A}_{t}}\leq 4\sigma\widetilde{\sigma}_{\max}\sqrt{\frac{2t(d+d_{h})\log(2Kt^{2}/\delta)}{p\lambda_{\min}\left(\mathbf{A}_{t}\right)}}.$
By Corollary 1, with $\epsilon\in(0,1)$ to be determined later, for $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log(2Kt^{2}/\delta)$, with probability at least $1-\delta/t^{2}$,
@@
Eq. 22 implies $(1-\epsilon)\mathbf{I}_{K}\preceq\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}$ and $(1-\epsilon)\mathbf{V}_{t}\preceq\mathbf{A}_{t}$. Thus,
@@
and Eq. 21 is bounded by,
$\begin{split}\left\|\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)\right\|_{\infty}\ &\leq\max_{i\in[K]}\left\|e_{i}^{\top}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1/2}\right\|_{2}\frac{4\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{\frac{2(d+d_{h})\log(2Kt^{2}/\delta)}{1-\epsilon}}\\ &\leq\max_{i\in[K]}\left\|e_{i}^{\top}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}\frac{4\sigma\widetilde{\sigma}_{\max}}{p(1-\epsilon)\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}.\end{split}$ (23)
By Eq. 22,
@@
$\leq\widetilde{\sigma}_{\max}\sqrt{t}\epsilon.$
Thus,
@@
Setting $\epsilon=K^{-1/2}\widetilde{\sigma}_{\min}\widetilde{\sigma}_{\max}^{-1}/8$ gives $21-\epsilon\geq 1/$ 2 and
@@
Now we conclude that for $t\geq 8^{3}K^{3}\widetilde{\sigma}^{-2}_{\min}\widetilde{\sigma}_{\max}^{2}(1-p)^{-2}\log(2Kt^{2}/\delta)$
@@
and by Lemma 3,
$\|\widehat{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\|_{\mathbf{V}_{t}}\leq 8\sigma\widetilde{\sigma}_{\max}\sqrt{\frac{2t(d+d_{h})\log(2Kt^{2}/\delta)}{p\lambda_{\min}\left(\mathbf{V}_{t}\right)}}=\frac{8\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}},$
@@
E.3 Proof of Theorem 3
Because the regret is bounded by $2$ and the number of rounds for the exploration phase is at most $|\mathcal{E}_{T}|\leq 8^{3}K^{3}(1-p)^{-2}\log(2KT^{2}/\delta)$.
$\mathrm{Reg}(T)$ $\leq 8^{3}K^{3}(1-p)^{-2}\log\frac{2KT^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]$
$=$ $8^{3}K^{3}(1-p)^{-2}\log\frac{2KT^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
$+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}.$
By Theorem 2, on the event $\{a_{t}=\widehat{a}_{t}\}$,
$\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]$ $=\widetilde{\boldsymbol{x}}_{a_{\star}}^{\top}{\boldsymbol{\mu}}_{\star}-\widetilde{\boldsymbol{x}}_{\widehat{a}_{t}}^{\top}{\boldsymbol{\mu}}_{\star}$
$\leq 2\max_{a\in[K]}\left|\widetilde{\mathbf{x}}_{a}^{\top}\left({\boldsymbol{\mu}}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{L}\right)\right|+\widetilde{\boldsymbol{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}-\widetilde{\boldsymbol{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}$
$\leq 2\max_{a\in[K]}\left|\widetilde{\mathbf{x}}_{a}^{\top}\left({\boldsymbol{\mu}}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{L}\right)\right|$
@@ E.3 Proof of Theorem 3
with probability at least $1-5\delta/t^{2}$ for each $t\in[T]\setminus\mathcal{E}_{T}$. Summing over $t$ gives,
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\leq\frac{16\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}}.$
By Lemma 5 and the Hoeffding bound, with probability at least $1-\delta$
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
$\leq 2\left[\sum_{t\in[T]}\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)-\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)+\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)\right]$
$\leq 4\sqrt{2T\log\frac{2}{\delta}}+\frac{4p\sqrt{T}}{K-1}+4\delta.$
Then with probability at least $1-6\delta$,
$\mathrm{Reg}(T)$ $\leq 8^{3}K^{3}(1-p)^{-2}\log\frac{2KT^{2}}{\delta}+\frac{4\sqrt{T}}{K-1}\frac{\log((T+1)^{2}/\delta)}{\log(1/p)}$
$\quad+4\sqrt{2T\log\frac{2}{\delta}}+4\delta+\frac{16\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}},$
@@
E.4 Proof of Theorem 4
Let $\widetilde{\mathbf{V}}_{t}:=\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}+\mathbf{I}_{K}$ and $\mathbf{V}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}+\mathbf{I}_{K}$. By definition of $\widehat{\boldsymbol{\mu}}_{t}^{R}$,

### v3

@@ E.2 Proof of Theorem˜2
$\leq\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\right)\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)\right\|_{\infty}+\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\boldsymbol{\mu}_{\star}\right)\mathbf{\widetilde{x}}_{a}\right\|_{\infty}.$ (20)
With probability at least $1-\delta/(\tau+1)^{2}$, the event $\mathcal{M}_{\tau}$ happens for all $\tau\geq 1$ and we obtain a pair of matching sample $\widetilde{a}_{\tau}$ and $a_{\tau}$. Under $\mathcal{M}_{\tau}$, the second term in Eq.˜20 is equal to,
@@
Because $\|\mathbf{v}\|_{\infty}=\max_{i\in[d]}|{\mathbf{e}_{i}^{\top}\mathbf{v}}|$ for any $\mathbf{v}\in\mathbb{R}^{d}$,
@@
$\leq\max_{a\in[K]}\sigma\sqrt{2\mathbf{e}_{a}^{\top}\mathbf{V}_{t}\mathbf{e}_{a}\log\frac{2Kt^{2}}{\delta}}$
@@
Now we turn to the first term in Eq.˜20. Define $\mathbf{A}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\mathbb{I}(\widetilde{a}_{\tau}=a)\phi_{a,\tau}^{-1}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$. Then the first term is rearranged as
@@
Because $\check{\boldsymbol{\mu}}_{t}^{L}$ is a minimizer of Eq.˜8, by Lemma 3 and Eq.˜21,
$\left\|\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right\|_{\mathbf{A}_{t}}\leq\frac{4\sigma\widetilde{\sigma}_{\max}}{p}\sqrt{\frac{2t(d+d_{h})\log(2Kt^{2}/\delta)}{\lambda_{\min}\left(\mathbf{A}_{t}\right)}}.$
By Corollary˜1, with $\epsilon\in(0,1)$ to be determined later, for $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log(2Kt^{2}/\delta)$, with probability at least $1-\delta/t^{2}$,
@@
Eq.˜23 implies $(1-\epsilon)\mathbf{I}_{K}\preceq\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}$ and $(1-\epsilon)\mathbf{V}_{t}\preceq\mathbf{A}_{t}$. Thus,
@@
and Eq.˜22 is bounded by,
$\left\|\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)\right\|_{\infty}\$$\leq\max_{i\in[K]}\left\|e_{i}^{\top}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1/2}\right\|_{2}\frac{4\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{\frac{2(d+d_{h})\log(2Kt^{2}/\delta)}{1-\epsilon}}$
$\leq\max_{i\in[K]}\left\|e_{i}^{\top}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}\frac{4\sigma\widetilde{\sigma}_{\max}}{p(1-\epsilon)\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}.$ (24)
For the first term in Eq.˜24, we have that
@@
$\leq\widetilde{\sigma}_{\max}\sqrt{t}\epsilon,$
where the last inequality holds due to Eq.˜23. Combining this result with Eq.˜24, we obtain that
@@
Setting $\epsilon=K^{-1/2}\widetilde{\sigma}_{\min}/8\widetilde{\sigma}_{\max}$ yields that $1-\epsilon\geq 1/2$ and that
@@
Now we conclude that for $t\geq 8^{3}K^{3}\widetilde{\sigma}^{-2}_{\min}\widetilde{\sigma}_{\max}^{2}(1-p)^{-2}\log(2Kt^{2}/\delta)$,
@@
by Lemma˜3 and taking a union bound on the both terms in Eq.˜20, with probability at least $1-2\delta/t^{2}$,
$\|\widehat{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\|_{\mathbf{V}_{t}}\leq\frac{8\sigma\widetilde{\sigma}_{\max}}{p}\sqrt{\frac{2t(d+d_{h})\log(2Kt^{2}/\delta)}{\lambda_{\min}\left(\mathbf{V}_{t}\right)}}=\frac{8\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}},$
@@
E.3 Proof of Theorem˜3
Since we have shown that the reward defined in Eq.˜2 is equivalent to its form obtained via the projection and augmentation strategy in Eq.˜7, we henceforth calculate the regret bound based on Eq.˜7. Under the assumption that the expected reward is bounded by 1 (Section˜3.2), the instantaneous regret $\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]$ is bounded above by $2$ for any $t\in[T]$, and that the number of rounds for the exploration phase satisfies $|\mathcal{E}_{T}|\leq 8^{3}K^{3}(1-p)^{-2}\log(2KT^{2}/\delta)$. Given these bounds, the cumulative regret is bounded as:
$\mathrm{Reg}(T)$ $=\sum_{t=1}^{T}(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}])$
$=\sum_{t\in\mathcal{E}_{T}}(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}])+\sum_{t\in[T]\setminus\mathcal{E}_{T}}(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}])$
$\leq 2\cdot 8^{3}K^{3}(1-p)^{-2}\log\frac{2KT^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}])$
$\begin{aligned} &=2\cdot 8^{3}K^{3}(1-p)^{-2}\log\frac{2KT^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}(a_{t}=\widehat{a}_{t})(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}])\right\}\\ &\quad+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}(a_{t}\neq\widehat{a}_{t})(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}])\right\}.\end{aligned}$ (25)
We first consider the second term of Eq.˜25. By Theorem˜2, on the event $\{a_{t}=\widehat{a}_{t}\}$,
$\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]$ $=\widetilde{\mathbf{x}}_{a_{\star}}^{\top}{\boldsymbol{\mu}}_{\star}-\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}{\boldsymbol{\mu}}_{\star}$
$=\widetilde{\mathbf{x}}_{a_{\star}}^{\top}\boldsymbol{\mu}_{\star}+\widetilde{\mathbf{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}-\widetilde{\mathbf{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}+\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}-\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}-\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}{\boldsymbol{\mu}}_{\star}$
$\leq\left|\widetilde{\mathbf{x}}_{a_{\star}}^{\top}(\boldsymbol{\mu}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{L})\right|+\left|{\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}(\boldsymbol{\mu}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{L})}\right|+\widetilde{\mathbf{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}-\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}$
$\leq 2\max_{a\in[K]}\left|\widetilde{\mathbf{x}}_{a}^{\top}({\boldsymbol{\mu}}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{L})\right|+\widetilde{\mathbf{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}-\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}$
$\leq 2\max_{a\in[K]}\left|\widetilde{\mathbf{x}}_{a}^{\top}({\boldsymbol{\mu}}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{L})\right|$
@@ E.3 Proof of Theorem˜3
with probability at least $1-2\delta/t^{2}$ for each $t\in[T]\setminus\mathcal{E}_{T}$. Summing over $t$ and applying a union bound, we obtain that, with probability at least $1-4\delta$,
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}(a_{t}=\widehat{a}_{t})\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\leq\frac{32\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}}.$ (26)
For the last term of Eq.˜25,
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}(a_{t}\neq\widehat{a}_{t})\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
$\leq 2\left[\sum_{t\in[T]}\mathbb{I}(a_{t}\neq\widehat{a}_{t})-\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)+\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)\right]\qquad(\because\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\leq 2)$
$\leq 2\sqrt{2T\log\frac{2}{\delta}}+\frac{4\sqrt{T}}{K-1}+4\delta,$ (27)
where the first term in Eq.˜27 holds with probability at least $1-\delta$ by Hoeffding’s inequality, and the remaining terms follow from Lemma˜5. Taking a union bound over Eq.˜26Eq.˜27, and $\mathcal{M}_{t}$, we obtain, with probability at least $1-6\delta$,
$\mathrm{Reg}(T)$ $\leq 2\cdot 8^{3}K^{3}(1-p)^{-2}\log\frac{2KT^{2}}{\delta}+\frac{4\sqrt{T}}{K-1}$
$\quad+2{\sqrt{2T\log\frac{2}{\delta}}}+4\delta+\frac{32\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}},$
@@
E.4 Proof of Theorem˜4
Let $\widetilde{\mathbf{V}}_{t}:=\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}+\mathbf{I}_{K}$ and $\mathbf{V}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}+\mathbf{I}_{K}$. By definition of $\widehat{\boldsymbol{\mu}}_{t}^{R}$ presented in Eq.˜12,